OpenStudy (anonymous):

Is this correct? (36x^2 - 49y^2) / (6x + 7y) = 36x^2 - 49y^2 = (6x - 7y)(6x - 7y) = (6x - 7y)(6x - 7y) / (6x + 7y) = 6x + 7y

OpenStudy (anonymous):

should be, if i calculated it right. :'3

OpenStudy (anonymous):

>.♥

OpenStudy (anonymous):

lol, ty n.n well hopefully I understood it right. :\ I'm not doing too hot in my math class.

OpenStudy (anonymous):

:P Me neither...idk why i'm in advanced.

OpenStudy (anonymous):

wait whats the would thing look like?

OpenStudy (anonymous):

.../ is the divide sign.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

X3

OpenStudy (anonymous):

brb phone call

OpenStudy (asnaseer):

you have a small mistake, it should be: 36x^2 - 49y^2 = (6x + 7y)(6x - 7y)

OpenStudy (asnaseer):

using:\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

ooh I see :3 ty for catching that n.n

OpenStudy (anonymous):

some are square roots

OpenStudy (asnaseer):

so your final answer needs to be changed

OpenStudy (anonymous):

so it would be 6x-7y ?

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

yay! ty Asnaseer ♥

OpenStudy (asnaseer):

yw